OpenStudy (anonymous):

Evaluate the rational function or expression for x=1 and x=-2. write undefined if appropriate. y=2/x

OpenStudy (anonymous):

please show work

OpenStudy (anonymous):

y = 2/x for x=1: y = 2/1 = 2 for x=-2: y = 2/(-2) = -1

OpenStudy (anonymous):

okay, thankyou!